# Closing the Gap on the Sample Complexity of 1-Identification

### v1

@@ B.2 Proof of Theorem 5
Following the sketch proof in section 2, we prove the lemmas one by one. The first step is to build a “constant probability lower bound”.
@@
Recall the notation $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, and the $\delta$-PAC assumption requires
@@
We prove the claim by contradiction, if $Pr_{\nu}(\tau<+\infty,\hat{a}\in\{a:\mu_{a}=\mu_{0}\})\geq 4\delta$ holds, we can find a $T_{0}$, such that
@@
We define $\nu^{\prime}(\epsilon)$ with $r_{a}^{\nu^{\prime}(\epsilon)}=\begin{cases}\mu_{0}-\epsilon&\mu_{a}=\mu_{0}\\ r_{a}^{\nu}&\text{else}\end{cases}$, where $\epsilon$ is an arbitrary positive number. We know $i^{*}(\nu^{\prime}(\epsilon))=[m]$ and we have
@@
The next step is to prove a lemma analyzing the pulling behavior.
@@
The main idea is to utilize Lemma 7. From the definition, we know $T_{j}(a)\geq 1$. For $n=1,2,\cdots,T_{j}(a)-1$, we have
@@
The third step is by the fact that $\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\}\subset\{N_{a}(\tau)+N_{j}(\tau)>n\}$ holds for all $n=0,1,\cdots,T_{j}(a)-1$. The last step is by the fact that $\frac{\Delta_{a,j}^{2}}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq 1$. For the same reason, we have
@@
Based on the Lemma 10, we can prove a lower bound connecting $\mathbb{E}N_{A}(\tau)$ and $\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)$, which is the following Lemma.
@@
We firstly prove a claim: Assume $\tau_{n+1}^{j}(a)<\tau$, for all $t$ such that $\tau_{n}+1^{j}(a)<t<\tau_{n+1}^{j}(a)+1$, we have $A_{t}\neq a$. We prove this claim by contradiction. Assume there exists $t$ such that $\tau_{n}^{j}(a)+1<t<\tau_{n+1}^{j}(a)+1$, $A_{t}=a$. Then, from the definition of $\tau_{n}^{j}(a)$, we know $N_{j}(\tau_{n}^{j}(a))+N_{a}(\tau_{n}^{j}(a))=n,N_{j}(\tau_{n}^{j}(a)+1)+N_{a}(\tau_{n}^{j}(a)+1)=n+1$. Since $t>\tau_{n}^{j}(a)+1$, which means $t-1\geq\tau_{n}^{j}(a)+1$, we can conclude
@@
However, $t\leq\tau_{n+1}^{j}(a)$ suggests that $N_{j}(t)+N_{a}(t)\leq N_{1}(\tau_{n+1}^{j}(a))+N_{a}(\tau_{n+1}^{j}(a))=n+1$, which leads to a contradiction. Then we can rewrite $\mathbb{E}_{\nu}N_{a}(\tau)$ as the following.
@@
$kl\Bigg(\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big),1-\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)\Bigg)<0.4$
where $kl(x,y)=x\log\frac{x}{y}+(1-y)\log\frac{1-y}{1-x}$. From Pinsker’s Inequality, we can further conclude
@@
The last line suggests that
@@
Given Lemma 11, we are able to take the optimal value of an optimization problem as the lower bound of $\mathbb{E}_{\nu}\tau$, which is Lemma 6.
@@
Thus, we can conclude
@@
Merge (18) and (19), notice that Lemma 9 asserts
@@
Then, we turn to show $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$ holds for $j\in[m],a\in[K]$. If $\mu_{a}\geq\mu_{j}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. If $\mu_{j}>\mu_{a}>\mu_{0}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}=\frac{1}{\Delta_{j,0}^{2}}>\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. If $\mu_{0}\geq\mu_{a}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}=\frac{1}{\Delta_{j,a}^{2}}>\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. Thus, we know $\mathbb{E}\tau$ must be lower bounded by the optimal value of the following optimization problem.
@@
Reformulate the above optimization problem, we can turn to solve
@@
$\geq$ $\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$
Before that, we prove a claim
@@
Define $\bar{L}_{j}(p)=\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p+\frac{p}{1+\log\frac{2}{p}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p$ and $p^{*}_{j}=\arg\min\limits_{1\geq p\geq 0}\bar{L}_{j}(p)$. Not hard to validate $\bar{L}_{j}(p)$ is a convex function. We prove the claim by discussing two cases.
@@
The second last line is by the definition of $\gamma^{0}$. The last line is by the fact that $2\sum_{s=1}^{m}\frac{1}{2s}<1+\log(m+1)$ for $m\geq 1$. This inequality contradicts with the fact that $\bar{L}_{j}(p^{*}_{j})<0$. Thus, we can conclude
@@ C.1 Rigorous Presentation of Algorithm
where $\lceil x\rceil:=\max\{x,1\}$
@@
where we consider $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$ as stacks that can add index to the stored item, and denote $A_{s}(\mathcal{H}^{\text{ee}}),X_{s}(\mathcal{H}^{\text{ee}})$ as the action and reward stored with index $s$ in $\mathcal{H}^{\text{ee}}$. Same idea applies to $A_{s}(\mathcal{H}^{\text{et}}),X_{s}(\mathcal{H}^{\text{et}})$. If $N_{a}^{\text{ee}}(\mathcal{H}^{ee})=0$ holds for some $a\in[K]$, we take $\text{UCB}^{\text{ee}}_{a}=+\infty,\text{LCB}^{\text{ee}}_{a}=$-$\infty$. Same idea applies to $N_{a}^{\text{et}}(\mathcal{H}^{et}),\text{UCB}^{\text{et}}_{a},\text{LCB}^{\text{et}}_{a}$. Given the above notation, we introduce algorithm 2, which is adapted from Li and Cheung (2025). The following analysis is very similar to the Section 4 and Appendix B in Li and Cheung (2025), but we still remain the full description and proof for completeness. In the following subsection, we discuss each part of the algorithm line by line and highlight our adaptation of history literature in Appendix C.1.5.
@@ 21            Take $\hat{a}_{k}=\textsf{None}$ and return None;
C.1.1 From Line 2 to 2: Input and Initialized Parameter
Algorithm 2 takes bracket $B$, threshold $\mu_{0}$, tolerance level $\delta$, arm number $K$, tunable parameter $C$ as input, and only pull arms in $B$. $\mu_{0},K$ are the same as what Algorithm 1 receives, while Algorithm 1 sets $\delta\leftarrow\frac{\delta}{\lceil\log_{2}K\rceil+1}$ when calling Algorithm 2. We take $C=1.01$ by default, but we can use any value greater than $1$.
Besides input parameter specified by the problem nature and Algorithm 1, 2 also initialize some parameters to manage the pulling process. Parameter $\{\delta_{k}=\frac{1}{3^{k}},\beta_{k}=2^{k},\alpha_{k}=5^{k}\}_{k=1}^{+\infty}$ are used to set up the tolerance level and budget during the pulling process. It is possible to specify other values for $\delta_{k},\beta_{k},\alpha_{k}$, see Appendix B.1 in Li and Cheung (2025). $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}},Q$ are used to store reward collected from pulling arms. As we have discussed above, $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$ are two stacks that can maintain the arrival order of reward of each arm $a\in B$. This requirement is to guarantee we are able to transfer the latest collected reward out of $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$. $Q$ is used to guarantee $\text{LCB}_{a}^{\text{ee}}$ cannot be greater than $\mu_{0}$ at the start of each phase, which will be discussed in section C.1.4.
@@ C.1.2 Line 2 to Line 2: Overall Idea of Algorithm 2
At the start of Algorithm 2, it splits the arm distributions into two parallel and independent copies, one for exploration period, and another for exploitation period. Correspondingly, reward collected in the exploration periods are stored in $\mathcal{H}^{\text{ee}}$. Reward collected in the exploration periods, which is from Line 2 to Line 2, are stored in $\mathcal{H}^{\text{et}}$.
After specifying the parameters through Line 2 and 2, Algorithm 2 splits the pulling process into multiple phases, noted by the For loop in Line 2. In each phase $k$, Algorithm 2 consists of two parts. The first part is the conditions from Line 2 to Line 2. The second part, which is called Exploration Period, ranges from Line 2 to 2. Exploration Period adopts UCB rule in Jamieson and Nowak (2014) to pull arms. These two parts are sequentially executed in the While Loop at Line 2. Though the condition of the While Loop is always True, but the conditions in the first part is able to guarantee that the Algorithm 2 never gets stuck in a infinity loop. See the discussion in section C.1.3.
At a phase $k$, the exploration period utilizes $\delta_{k}$ as the tolerance level to apply UCB rule of pulling arms, while the exploitation period utilizes $\frac{\delta}{\alpha_{k}}$ as the tolerance level to calculate Upper and Lower Confidence Bounds and only pull a single arm. Once the tolerance level of both periods are below $\frac{\delta}{3}$ and $\hat{a}$ takes value in $K]\cup\{\textsf{None}\}$, Algorithm 2 will terminate. What’s more, Algorithm 2 also sets up a maximum pulling times for both periods. For the exploration period at phase $k$, the maximum pulling times for each arm $a\in B$ is $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$. For the exploitation period at phase $k$, the maximum pulling times of the single arm is $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$. Section C.1.3 and C.1.4 provide more details.

### v2

@@ B.2 Proof of Theorem 5
Following the sketch proof in section 2, we prove the following lemmas one by one. The first step is to build a “probabilistic constant lower bound”. Without further description, we only consider symmetric algorithms in this section.
@@
Recall the notation $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$. The $\delta$-PAC assumption implies
@@
Before the proof of the Lemma 10, we first show a claim
@@
This claim is to show $\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])\geq 1-\Theta(\delta)$. If $\{a:\mu_{a}=\mu_{0}\}=\emptyset$, (15) is immediate. We therefore assume $\{a:\mu_{a}=\mu_{0}\}\neq\emptyset$ and argue by contradiction. If $Pr_{\nu}(\tau<+\infty,\hat{a}\in\{a:\mu_{a}=\mu_{0}\})\geq 4\delta$ holds, then there exists a (possibly $\delta$-dependent) constant $T_{0}(\delta)$, such that
@@
We define $\nu^{\prime}(\epsilon)$ with $r_{a}^{\nu^{\prime}(\epsilon)}=\begin{cases}\mu_{0}-\epsilon&\mu_{a}=\mu_{0}\\ r_{a}^{\nu}&\text{else}\end{cases}$, where $\epsilon$ is an arbitrary positive number. We know $i^{*}(\nu^{\prime}(\epsilon))=[m]$. By $\delta$-PAC assumption,
@@
We next quantify the algorithm’s inability to distinguish two arms when the total arm pulls of the pair is small.
@@
Dividing $\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$ on both sides, the conclusion is equivalent to
$\text{kl}\Big(\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a|\tau_{n}^{j}(a)<\tau),1-\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a|\tau_{n}^{j}(a)<\tau)\Big)\leq 0.4$
where $\text{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y}$. The reason is $\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=j)+\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=a)=\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$.
@@
The main idea is to utilize Lemma 8. Take $\Delta=\Delta_{a,j}$. From the definition, we know $T_{j}(a)\geq 1$. For $n=1,2,\cdots,T_{j}(a)-1$, we have
@@
The third line is by the fact that $\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\}\subset\{N_{a}(\tau)+N_{j}(\tau)>n\}$ holds for all $n=0,1,\cdots,T_{j}(a)-1$. The last step is by the fact that $\frac{\Delta_{a,j}^{2}}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq 1$ and
$\frac{\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}}{1+\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}}\leq 1.$
Similarly, we have
@@
The next lemma converts Lemma 11 into a conclusion relating $\mathbb{E}_{\nu}N_{a}(\tau)$ to the probability of correctly recommending a given qualified arm with sufficiently many samples.
@@
We first claim that if $\tau_{n+1}^{j}(a)<\tau$, for all $t$ such that $\tau_{n+1}^{j}(a)<t<\tau_{n+1}^{j}(a)+1$, we have $A_{t}\neq a$. We prove this claim by contradiction. Assume there exists $t$ such that $\tau_{n}^{j}(a)+1<t<\tau_{n+1}^{j}(a)+1$, $A_{t}=a$. Then, from the definition of $\tau_{n}^{j}(a)$, we know $N_{j}(\tau_{n}^{j}(a))+N_{a}(\tau_{n}^{j}(a))=n,N_{j}(\tau_{n}^{j}(a)+1)+N_{a}(\tau_{n}^{j}(a)+1)=n+1$. Since $t>\tau_{n}^{j}(a)+1$, which means $t-1\geq\tau_{n}^{j}(a)+1$, we can conclude
@@
However, $t\leq\tau_{n+1}^{j}(a)$ suggests that $N_{j}(t)+N_{a}(t)\leq N_{1}(\tau_{n+1}^{j}(a))+N_{a}(\tau_{n+1}^{j}(a))=n+1$, which leads to a contradiction. The claim follows.
Using the claim, we can rewrite
@@
$\text{kl}\Bigg(\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big),1-\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)\Bigg)<0.4$
where $\text{kl}(x,y)=x\log\frac{x}{y}+(1-y)\log\frac{1-y}{1-x}$. From Pinsker’s Inequality, we can further conclude
@@
Combine the last line and (17), we get
@@
Finally, we use Lemma 12 to lower bound $\mathbb{E}_{\nu}\tau$ by the optimal value of an optimization problem, as stated in Lemma 6.
@@
The third step is by the Lemma 12. Since the last line holds for any $j\in[m]$,
@@
Merge (18), (19), and notice that Lemma 10 asserts
@@
where for any symmetric and $\delta$-PAC algorithm, $\Big\{p_{j}=\Pr_{\nu}\Big(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Big)\Big\}_{j=1}^{m}$ is always a feasible solution.
The next step is to show $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$ holds for $j\in[m],a\in[K]$. If $\mu_{a}\geq\mu_{j}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. If $\mu_{j}>\mu_{a}>\mu_{0}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}=\frac{1}{\Delta_{j,0}^{2}}>\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. If $\mu_{0}\geq\mu_{a}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}=\frac{1}{\Delta_{j,a}^{2}}>\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. Thus, we know $\mathbb{E}\tau$ must be lower bounded by the optimal value of the following optimization problem.
@@
Reformulate the above optimization problem, we know the optimal value of (2)
@@
$\geq$ $\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$ (20)
We start the proof of (20) by claiming
@@
holds for all $j\in[m]$. Define $\bar{L}_{j}(p)=\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p+\frac{p}{1+\log\frac{2}{p}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p$ and $p^{*}_{j}=\arg\min\limits_{1\geq p\geq 0}\bar{L}_{j}(p)$. Not hard to validate $\bar{L}_{j}(p)$ is a convex function for $p\in[0,1]$, and there are two cases.
@@
The second last line is by the definition of $\gamma^{0}$. The last line is by the fact that $2\sum_{s=1}^{m}\frac{1}{2s}<1+\log(m+1)$ for $m\geq 1$. This last line contradicts with the fact that $\bar{L}_{j}(p^{*}_{j})<0$. Thus, we can conclude $p^{*}_{j}<\frac{1}{100m}$, further
@@ C.1 Rigorous Presentation of Algorithm
where $\lceil x\rceil:=\max\{x,1\}$.
@@
We treat $\mathcal{H}^{\mathrm{ee}}$ and $\mathcal{H}^{\mathrm{et}}$ as stacks with indexed entries: $A_{s}(\mathcal{H}^{\mathrm{ee}})$ and $X_{s}(\mathcal{H}^{\mathrm{ee}})$ denote, respectively, the action and reward stored at index $s$ in $\mathcal{H}^{\mathrm{ee}}$, and analogously for $\mathcal{H}^{\mathrm{et}}$. When $N_{a}^{\mathrm{ee}}(\mathcal{H}^{\mathrm{ee}})=0$ for some $a\in[K]$, we set $\mathrm{UCB}^{\mathrm{ee}}_{a}=+\infty$ and $\mathrm{LCB}^{\mathrm{ee}}_{a}=$-$\infty$; the same convention applies to $\mathrm{UCB}^{\mathrm{et}}_{a}$ and $\mathrm{LCB}^{\mathrm{et}}_{a}$ when $N_{a}^{\mathrm{et}}(\mathcal{H}^{\mathrm{et}})=0$.
With this notation in place, we present Algorithm 2, adapted from Li and Cheung [2025]. Our analysis follows closely Section 4 and Appendix B of Li and Cheung [2025]; nevertheless, we provide the full algorithmic description and proofs for completeness. In the subsequent subsections, we discuss the algorithm line by line and summarize our modifications relative to the prior literature in Appendix C.1.5.
@@ 21            Take $\hat{a}_{k}=\textsf{None}$ and return None;
C.1.1 From Line 2 to 2: Input and Initialize Parameter
Algorithm 2 takes a bracket $B$, a threshold $\mu_{0}$, a tolerance level $\delta$, the number of arms $K$, tunable parameter $C$ as input, and only pull arms in $B$. The parameters $\mu_{0}$ and $K$ coincide with those used by Algorithm 1. When Algorithm 1 calls Algorithm 2, it sets $\delta\leftarrow\delta/(\lceil\log_{2}K\rceil+1)$. We use $C=1.01$ by default, although any value strictly larger than $1$ is admissible.
In addition to the problem parameters, 2 also initialize some parameters to manage the pulling process. Parameter $\{\delta_{k}=\frac{1}{3^{k}},\beta_{k}=2^{k},\alpha_{k}=5^{k}\}_{k=1}^{+\infty}$ are used to set up the tolerance level and budget during the pulling process. It is possible to specify other values for $\delta_{k},\beta_{k},\alpha_{k}$, see Appendix B.1 in Li and Cheung [2025]. The containers $\mathcal{H}^{\mathrm{ee}}$, $\mathcal{H}^{\mathrm{et}}$, and $Q$ store samples collected throughout the process. In particular, $\mathcal{H}^{\mathrm{ee}}$ and $\mathcal{H}^{\mathrm{et}}$ are ordered stacks that preserve the arrival order of samples for each arm in $B$, which allows us to remove and transfer the most recently collected sample when needed. The role of $Q$ is to ensure that $\mathrm{LCB}^{\mathrm{ee}}_{a}$ cannot be greater than $\mu_{0}$ at the beginning of each phase; see Section C.1.4.
@@ C.1.2 Line 2 to Line 2: Overall Idea of Algorithm 2
At the start of Algorithm 2, it splits the arm distributions into two parallel and independent copies. One copy is for exploration period, and another copy is for exploitation period. Correspondingly, reward collected in the exploration periods are stored in $\mathcal{H}^{\text{ee}}$. Reward collected in the exploration periods, which is from Line 2 to Line 2, are stored in $\mathcal{H}^{\text{et}}$.
After specifying the parameters through Line 2 and 2, Algorithm 2 splits the pulling process into multiple phases, noted by the For loop in Line 2. In each phase $k$, Algorithm 2 consists of two parts. Each phase consists of two components executed within the while-loop in Line 2:
  • a sequence of stopping/transition checks (Lines 22), and
  • an exploration period (Lines 22) that applies a UCB rule (as in Jamieson and Nowak [2014]) to select the next arm to sample.
Although the loop condition is True, the stopping/transition checks ensure that the algorithm cannot run indefinitely; see Section C.1.3.
Within phase $k$, exploration uses tolerance level $\delta_{k}$ to form UCB/LCB bounds, while exploitation uses tolerance level $\delta/\alpha_{k}$ and samples only a single candidate arm. Once $\hat{a}$ takes value in $K]\cup\{\textsf{None}\}$, Algorithm 2 will terminate. Moreover, each phase enforces explicit sampling budgets: during exploration, each arm $a\in B$ can be sampled at most $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ times, while during exploitation the candidate arm is sampled at most $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$ times. Further details appear in Sections C.1.3 and C.1.4.